### v1

@@ Abstract
We formulate the Knapsack Contracts problem—a strategic version of the classic Stochastic Knapsack problem, which builds upon the inherent randomness shared by stochastic optimization and contract design. In this problem, the principal incentivizes agents to perform jobs with stochastic processing times, the realization of which depends on the agents’ efforts.
Algorithmically, we show that Knapsack Contracts can be viewed as Stochastic Knapsack with costs and multi-choice, features that introduce significant new challenges. We identify a crucial and economically meaningful parameter—the Return on Investment (ROI) value. We show that the Inverse of ROI (or IOR for short) precisely characterizes the extent to which the approximation guarantees for Stochastic Knapsack extend to its strategic counterpart.
For IOR of $\alpha$, we develop an algorithm that finds an $O(\alpha)$-approximation policy that does not rely on adaptivity. We establish matching $\Omega(\alpha)$ lower bounds, both on the adaptivity gap, and on what can be achieved without full distributional knowledge of the processing times. Taken together, our results show that IOR is fundamental to understanding the complexity and approximability of Knapsack Contracts, and bounding it is both necessary and sufficient for achieving non-trivial approximation guarantees. Our results highlight the computational challenges arising when stochasticity in optimization problems is controlled by strategic effort.
@@ 1 Introduction
The study of how incentives reshape classic algorithmic problems has been at the heart of algorithmic game theory research since its inception. In their seminal work, Nisan and Ronen [24] famously conjectured that adding incentives for truthfulness to the optimization problem of scheduling would make the problem extremely hard to approximate. Two decades later, intensive research has culminated in a proof of this conjecture [9]. Many additional works explore the intersection of algorithm design with truthfulness constraints [27, 25, 10, 6, 8, 22, 23].
Over recent years, the algorithmic game theory community has begun to explore incentive constraints beyond truthfulness. A recent focus is on an agent’s earnestness in exerting effort to complete a task, where the task is delegated to the agent by a principal. The main economic tool for incentivizing earnest effort is performance-based payment schemes, a.k.a. contracts.
Contracts are designed to create incentives that counteract moral hazard: When an agent exerts effort to carry out a job on behalf of a principal, this leads to a stochastic outcome that only partially reflects the agent’s effort. This stochasticity opens the possibility for the agent to claim they worked hard despite exerting a low level of effort, by blaming poor outcomes on “bad luck”. This phenomenon is known as moral hazard. A contract uses outcome-based payments and distributional knowledge to incentivize earnest effort, thus mitigating moral hazard. The goal of this work is to initiate the study of the intersection of algorithm design with contractual incentives for earnest effort.
1.1 Problem Formulation
To achieve our goal, we seek an algorithmic optimization problem with a natural connection to contractual incentives—similar to the connection between scheduling and truthfulness [24]. Our starting point is the observation that the field of contract design has a natural affinity to stochastic optimization, due to the nondeterminism inherent to both. Building upon this connection, we formulate a model of Knapsack Contracts, in which jobs that need to fit into a given time frame are delegated to agents. In our model, the agents have costs for carrying out the jobs, and the completion times (a.k.a. job lengths or sizes) are a stochastic outcome of the level of effort. The latter makes our model a generalization of Stochastic Knapsack [11]. We show that the agents’ costs lead to incentive constraints that make the problem significantly harder than classic Stochastic Knapsack. Since Stochastic Knapsack has a central place in stochastic optimization [3, 19, 20, 18, 28, 5, 2, 21], we view its study through the prism of contract design as a potential gateway to deeper exploration of the intersection between the two areas.
Example 1.
A PI (principal) wishes to run a set of experiments on the cloud before the deadline of a top AI conference. Each experimental task has a fixed value for the paper, which does not depend on how fast the cloud providers (agents) finish running the experiments. The PI does not benefit from experiments that finish running after the submission deadline. Despite this, she still needs to pay the providers for running them. The PI must spend substantial funds from her research grant to incentivize the agents to use powerful but costly resources when processing the experiments, resulting in shorter running times. This leads to an interesting tension between the PI’s benefit from an experiment, and how much they are willing to pay to incentivize an agent to run it quickly. The problem of optimizing the experiments thus has a mixed-sign objective.
Generalizing Example 1, we consider a single principal and multiple agents with jobs that must be executed sequentially. Each agent $i$ is associated with a fixed known value $v_{i}$ to the principal for completing their job before the deadline expires. Moreover, each agent can choose one of $m$ costly actions, and the chosen action $j\in[m]$ determines a distribution $F_{i}^{j}$ over the possible completion times of the job. The principal cannot directly observe the action taken by an agent, and can only inspect the completion times. To incentivize agent $i$, the principal designs a contract $t_{i}$ that maps the completion time $s_{i}$ to a monetary transfer. Each agent $i$ wants to maximize their own utility: the transfer received minus the cost $c_{i}^{j}$ exerted by them when taking action $j$. The principal wants to maximize the profit, which is the sum of values over all jobs completed before the deadline, minus the sum of transfers made to the agents (i.e., a mixed-sign objective). See Section 2 for a formal problem description.
Relation to vanilla contract design.
Usually in contract settings, the value created for the principal by the agent’s effort is stochastic. Our model captures settings in which the principal’s value for the job delegated to the agent is fixed, but the time it takes the agent to complete the job is a random variable, and if the job doesn’t finish on time then the resulting value is zero. This seemingly simple twist on the standard principal-agent model leads to new challenges.
Relation to Stochastic Knapsack.
In the standard Stochastic Knapsack problem, there are $n$ jobs to execute, each with a fixed value but stochastic size. A policy-maker executes jobs sequentially, gaining the value of any job that terminates before the deadline. This process ends once a job terminates after the deadline. The optimal policy for this problem can be highly complex, adaptive, and intractable. Yet, Dean et al. [11] show that there exists a simple non-adaptive policy that approximates, up to a constant factor, the reward of the optimal adaptive policy.
In contrast, in our model, even if a job ends after the deadline, the principal must pay the agent according to the contract, for processing it. As we show in Section 2.2, our model corresponds to a version of Stochastic Knapsack with job costs and multi-choice, and we show this makes a substantial difference relative to [11]. A first indication of the difference arises from the linear programming approach of [11], which they use to establish their main result. In their setting, the integrality gap of the LP is bounded; however, in Appendix A we demonstrate that the LP approach can grossly overshoot the true optimum in our setting. In fact, as we shall see, in our model there is no constant-approximation non-adaptive policy; any constant-approximation algorithm must adaptively choose the next job to execute. Intuitively, adaptivity is necessary to avoid the danger of incurring a very high cost, so high it overshadows any value the algorithm may obtain.
1.2 Our Results and Techniques
@@
The example in Appendix A is based on a “hard instance” of Stochastic Knapsack with job costs. We show this instance (and in fact any other hard instance) has the following property: For some of the jobs, their cost almost “covers” their value. Thus, in order to incentivize such jobs, the principal has to pay the agent their cost, and therefore, they are left with a small return on investment (ROI). I.e., the ratio between the obtained profit and the transfer made is small. Such jobs are intuitively problematic—if their cost must be reimbursed by the principal, but their value is not gained since they finish after the deadline, the loss can be substantial. We essentially show that these are the only problematic jobs.
In order to parameterize this, we inspect the Inverse of the ROI (or IOR) of a job (as the ROI becomes small, the IOR grows). We introduce an economically-meaningful parameter $\alpha$, which represents the worst IOR among all jobs—see Definition 2.6. Our main finding is that IOR is precisely the parameter that determines the loss in approximation due to moving from the classic Stochastic Knapsack problem, to a contract design problem subject to incentive constraints:
  • We show that the principal can efficiently and non-adaptively compute a set of contracts, which give an $O(\alpha)$-approximation to the profit of the (intractable and highly adaptive) optimal policy of the principal. See Theorem 3.1.
  • We show the existence of an instance such that any non-adaptive policy of the principal is bounded away by a factor of $\Omega(\alpha)$ in terms of profit from the optimal adaptive policy. This establishes a tight $\Theta(\alpha)$ adaptivity gap for the Knapsack Contracts problem. See Theorem 4.1.
  • We establish an information gap: Consider any policy (adaptive or not) that has limited information about the jobs’ size distributions in the form of finitely many moments, as well as access to the probability of knapsack overflow at every step. No such policy can guarantee an approximation factor better than $O(\alpha)$ with respect to the profit of the optimal policy with full access to the size distributions. See Theorem 5.2.
We now describe our results in more detail as well as the techniques that enable them.
Positive results.
In Section 3, we provide an efficient algorithm that finds a non-adaptive set of contracts (see Section 2 for adaptivity notions), which give an $O(\alpha)$-approximation to the optimal profit the principal can obtain, where $\alpha$ is the IOR parameter (Theorem 3.1). Our algorithm needs to choose in advance the order in which agents are to execute their jobs and which effort level to incentivize each agent. Each effort level induces a different expected transfer needed in order to incentivize it. As a first step, in Lemma 2.1, we reduce the Knapsack Contract problem to a variant of Stochastic Knapsack that we term the Multi-Choice Stochastic Knapsack Problem with Costs (MSKC). In this purely algorithmic problem, each job is assigned one of multiple combinations of cost and size distribution. When a policy attempts to execute a job, it has to pay the cost up front.
@@
Our algorithm formulates an LP (see Eq. (2)) for the MSKC problem, where the weight of each job $i$ and effort level $j$ corresponds to the profit of executing job $i$ at effort level $j$, and the size is the expected size of the job at effort level $j$. In Dean et al. [11], an LP formulation of the problem gives an upper bound on the reward of the optimal adaptive policy, which is approximated by the reward of the best non-adaptive policy. In our case, the LP is still an upper bound on the optimal profit of an adaptive policy, but this upper bound can be a factor $\alpha$ away (see Appendix A for an example). We show that there exists a non-adaptive set of contracts that yields an $O(\alpha)$-approximation to the LP’s value.
The way in which we construct the set of contracts is the following. We first efficiently find a well-structured solution to the LP where each job is taken integrally at up to one effort level, except for at most a single job that can be taken as a convex combination of two effort levels. This is done by taking an arbitrary solution of the LP, and then shifting weights along the Pareto curves of jobs in order to iteratively move jobs to be taken integrally, while keeping the same solution value. Once we obtain the well-structured solution, we either use a standard bang-per-buck ordering of the jobs selected, or take a single job if its expected profit is larger than a predefined threshold. This threshold is set in a way that will (i) prevent a large loss for the bang-per-buck order in case the IOR parameter $\alpha$ has a large value; and (ii) ensure a large profit when there’s a single job that gives such profit, but might have a low probability of being selected when jobs are executed in the bang-per-buck order.
Negative results.
In Section 4 and Section 5 we give further evidence for the role that IOR has in Knapsack Contracts. In Section 4 we inspect the adaptivity gap, considering three types of policies:
  1. 1.
    Fully adaptive policies that can choose the next job to execute (if any) and at which effort level based on the entire history.
  2. 2.
    Stopping-time adaptive policies that chooses a fixed set of contracts (job/effort level pairs) up front at a fixed order, but stops once any future sequence of jobs has a negative expected profit. These types of policies avoid a large loss that the principal can incur due to the transfer component of the contract problem and are used in the algorithm of Section 6.
  3. 3.
    Non-adaptive policies that chooses a fixed set of contracts at a fixed order, and the process terminates either when the last job in the order has been executed, or when there’s a job that is finished after the deadline.
Dean et al. [11] show that in the setting without costs, the reward of the optimal non-adaptive policy is within a factor 4 of the optimal adaptive policy. We show that when introducing costs, the ratio between these can be unbounded. Namely, for every IOR parameter $\alpha>1$, we show:
  • An instance where the ratio between the profit of stopping-time adaptive policies and non-adaptive policies is $\Omega(\alpha)$ (Section 4.1).
  • An instance where the ratio between the profit of fully adaptive policies and stopping-time adaptive policies is $\Omega(\alpha^{1/3})$ (Section 4.2).
Our instances are ones where each job has only a single effort level, demonstrating that the cost aspect by itself leads to a significant change relative to the setting without cost.
Information barrier.
In Section 5 we inspect how much information is needed to get an $O(\alpha)$ approximation in the purely algorithmic Stochastic Knapsack problem where there is an inherent cost for executing jobs (this corresponds to having a single effort level for each agent). In the purely algorithmic setting, our algorithm from Section 3 only uses first moment information (the expectation) of jobs’ sizes, along with the probability each job’s size will exceed the deadline, and obtains an $O(\alpha)$ approximation. We show that this information does not suffice if one wants to obtain a better approximation. Namely, for every finite $k$, there exists an instance where any (possibly adaptive) policy that can only access the $k$ moments of the size distribution, along with the probability that each job will exceed the deadline, is bounded away by a factor of $\Omega(\alpha)$ from the optimal policy (Theorem 5.2).
Bounded-size jobs.
In Section 6, we consider the case where the IOR parameter can be arbitrarily large, but each job’s probability to take more than $1-\delta$ of the length of the time budget is small for $\delta\in(0,1)$. We give an efficient algorithm that finds a stopping-time adaptive set of contracts, which guarantees an $O(1/\delta)$ approximation to the optimal profit of a fully-adaptive set of contracts.
In order to obtain this result, we first consider a variant of the problem where, when a job finishes after the deadline, the principal still obtains its reward, and then the process stops. For this variant, the algorithm from Section 3 gives a constant approximation to the LP value. We then consider running this algorithm while imagining the deadline is $\delta$-fraction of the actual deadline. We show that in this case, having bounded-size jobs is equivalent to obtaining the reward of the first job that finishes after the deadline. Moreover, halving the deadline only causes a factor $\delta$ loss in the objective. Note that (moderate) adaptivity is required for this variant as well, as non-adaptive policies lose a factor that grows with $\alpha$ (see Appendix B).
Ordered-Adaptive Contracts.
Motivated by real-world examples where jobs must be executed in a given order, in Appendix D we study the model where the order is given, but for each job in the order, the principal can decide whether to execute it, and if so, at which effort level. In Theorem D.1 we give a resource augmentation-type result: given a time budget larger by a factor of $1+\epsilon$ from the original time budget, one can compute an adaptive policy efficiently (in polynomial time in the number of jobs, effort levels, and $1/\epsilon$), which obtains a reward larger than the optimal adaptive policy for the original time budget. The policy, similar to [11, 5], defines a dynamic programming solution on carefully discretized rounded-down processing times. This result hints that our problem boils down to ordering the jobs; i.e., one may hope to find a nearly-optimal adaptive policy for our non-ordered model if one can efficiently find a near-optimal ordering of the jobs.
1.3 Related Work
Classic contract theory.
Classic contract theory has a long-standing tradition in economics, focusing on incentive design under asymmetric information. A result by Carroll [7] shows that linear contracts are often robust across a range of environments, providing a foundation for the study of simpler contract forms. These classical models emphasize moral hazard, adverse selection, and trade-offs between risk and incentives.
Algorithmic contract theory.
In recent years, a series of studies has emerged that study contracts through the lens of the theory of computation, initiated by [12, 4]. The authors of [12] raise concerns regarding the complexity and unintuitiveness of optimal contracts, while exploring the approximation ratios of simpler contract forms, such as linear contracts. The multi-agent setting was initially explored by Babaioff et al. [4]. A recent survey paper [13] provides an overview of developments in algorithmic contract theory.

### v2

@@ Abstract
We revisit the Stochastic Knapsack problem, where a policy-maker chooses an execution order for jobs with fixed values and stochastic running-times, aiming to maximize the value completed by a deadline. Dean et al. (FOCS’04) show that simple non-adaptive policies can approximate the (highly adaptive) optimum, initiating the study of adaptivity gaps.
We introduce an economically motivated generalization in which each job also carries an execution cost. This uncovers new applications, most notably a new and natural variant of contract design: jobs are processed by agents who choose among effort levels that induce different processing-time distributions, and the principal must decide which jobs to run and what payments induce the intended effort.
With costs, the objective becomes mixed-sign: value from completed jobs must be balanced against costs of execution, and running a job that misses the deadline can create negative utility. This changes the algorithmic picture, and the adaptivity gap is no longer constant. We give an economic explanation: the performance of non-adaptive policies is governed by jobs’ return on investment (ROI) — utility over cost — which can be arbitrarily small.
We introduce a hierarchy of increasingly adaptive policies, trading off simplicity and adaptivity. We prove near-tight guarantees across the hierarchy, showing that with costs the adaptivity gap is $\Theta(\alpha)$, where $1/\alpha$ is the ROI. Higher in the hierarchy, we identify an efficiently computable policy with limited adaptivity that is approximately-optimal. Analogous to the centrality of ROI in economics, we believe our ROI-based, simple-vs-optimal approach to adaptivity may be useful for additional stochastic optimization and online problems with mixed-sign objectives.
@@ 1 Introduction
In many economic decision-making settings, actions have stochastic outcomes and are associated not only with value but also with inherent costs. While most of the existing literature focuses on actions that generate value alone, in this paper we study how incorporating costs can significantly change the algorithmic design of effective solution methods. Understanding how costs affect decision-making scenarios is a central challenge in designing practical, near-optimal policies in many practical settings.
Stochastic Knapsack and Adaptivity Gaps.
In this work, we consider one of the classic stochastic decision problems studied in the seminal paper of Dean et al. [10]: the Stochastic Knapsack problem. In this problem, a policy-maker designs a policy that sequentially executes jobs, each associated with a known value and an uncertain processing time. The policy-maker faces a fixed deadline and obtains value only from jobs that complete before this deadline; any job finishing after the deadline yields zero value. While processing times are uncertain, the policy-maker is assumed to know their distributions, and seeks a policy that maximizes the expected total value obtained before the deadline.
A key insight of Dean et al. [10] is the trade-off between policy simplicity and optimality, formalized through the notion of adaptivity gap. The optimal policy may adaptively choose the next job to process based on the realized processing times of previously executed jobs. Such adaptivity can be extremely complex: the optimal policy can be represented as a decision tree, whose size might be exponential in the number of jobs.
Motivated by this potential complexity, Dean et al. [10] study simple non-adaptive policies, in which jobs are ordered in advance and executed sequentially until the deadline, collecting rewards from those that complete in time. Despite their simplicity, they show that such policies achieve a constant-factor approximation of the optimal adaptive policy, implying that the adaptivity gap is bounded by a constant. While the results of Dean et al. [10] demonstrate the surprising power of simple non-adaptive policies, they rely on the assumption that jobs can be executed without incurring costs.
Costly Job Execution and Multiple Execution Modes.
In many fundamental applications, assuming job execution is cost-free is unrealistic. We therefore study a setting with costly job execution, motivated by applications such as cloud computing or contracting with workers, where executing a job requires explicit compensation and thus incurs a cost.
Moreover, each job may be executed at multiple levels of effort, where higher effort corresponds to higher cost, and different efforts induce different distributions over processing times.
For example, the choice of the machine used to process a job, or of the specialist hired to perform a task, affects both the execution cost and the processing-time distribution.
Although these extensions may appear modest, they fundamentally change the role of adaptivity. Notably, we show that when execution costs are present, the adaptivity gap of non-adaptive policies can be unbounded. Nevertheless, we identify a natural and economically meaningful parameter of the setting—capturing a notion of return on investment (ROI)—that governs the adaptivity gap in the presence of costs. We then show that a limited amount of carefully-applied adaptivity suffices to obtain policies whose performance is within a constant factor of the optimal adaptive policy.
@@
We notice that “hard instances” of our problem have the following structural property: For some of the jobs, their cost almost eclipses their value. Thus, when executing such jobs, they provide a small return on investment (ROI). More formally, the ratio between the obtained utility (value minus cost) from running the job and the incurred cost is small.
Such jobs are intuitively problematic in the knapsack context — their execution adds to the policy-maker’s cost, but their value might be lost if they finish after the deadline, in which case the policy-maker’s loss can be substantial. We essentially show that these are the only problematic jobs.
Hierarchy of Adaptivity.
Dean et al. [10] studied the Stochastic Knapsack problem by considering two extremes:
  • Non-adaptive policies. These policies choose an ordered set of jobs. The set is then processed in order until either all jobs are completed or the deadline is met, whichever occurs first.
  • Fully-adaptive policies. This is the most general policy class: there is no predetermined job order. The policy chooses which job to run next based on the observed runtime history.
When costs are introduced, simple non-adaptive policies no longer guarantee a constant approximation to the value of the optimal fully-adaptive policies. We explore the landscape through the simplicity-optimality lens, defining and analyzing the following intermediate classes of policies:
  • Stopping-adaptive policies. Policies in this class choose an ordered set of jobs, but stop executing once, given the time left until the deadline, any future sequence of jobs results in a negative expected utility.
  • Skipping-adaptive policies. These policies also commit to an ordered list, but unlike stopping-adaptive policies, they may adaptively skip some of the listed jobs.
See Section 2.2 for formal definitions and how the decision tree policy representation size decreases along the hierarchy. In the appendix we also consider non-stopping policies, which are non-adaptive policies that never halt and execute all jobs in the ordered set even after the deadline has passed.
When there are multiple execution modes in addition to costs, the above classes of policies extend naturally by ordering pairs of (job, execution mode) rather than just ordering jobs.
Motivating Example.
To illustrate the differences among these policy types, consider the following motivating example, in which a contractor hires subcontractors to perform jobs under a time constraint. The contractor’s level of commitment and flexibility varies across policy classes.
Under a fully-adaptive policy, the contractor never commits to an execution order in advance. After each observed runtime, the contractor may choose any remaining job and execution mode next, redesigning the plan on the fly as a function of the realized history and the remaining time. Such a high degree of adaptivity, however, can significantly complicate the system, both for the contractor, who must continuously re-optimize decisions, and for the subcontractors, who face uncertainty regarding their scheduling times and effort level.
We therefore turn to simpler policy classes that impose an underlying fixed execution order on the jobs, specified by a priority list of job–execution-level pairs determined in advance. Under non-adaptive policies, once this priority list is fixed, the contractor is obliged to continue attempting jobs in order whenever any time remains. In this setting, a job that is highly unlikely to finish on time will still get executed,
incurring its execution cost without obtaining its value. As we show, this behavior can be highly inefficient, particularly when the return on investment of the final job is low.
To mitigate this issue, stopping-adaptive policies introduce a limited form of adaptivity by allowing the contractor to stop execution once future jobs cannot yield positive utility. Unfortunately, as we demonstrate, this additional adaptivity is insufficient to guarantee a constant-factor approximation.
Finally, we consider a class of skipping-adaptive policies that augments a fixed priority order with job-specific time thresholds. A job is executed upon reaching its position in the order only if the remaining time before the deadline is at least its threshold; otherwise, it is skipped. We show this rather simple policy class is expressive enough to guarantee a constant-factor approximation.
Application to Contract Design — Knapsack Contracts.
Our initial motivation in studying the stochastic knapsack with costs problem was to study a variant of the principal-agent model, which we term Knapsack Contracts. In this model, the value of each executed task is fixed, but the duration of the task may depend on the effort level of the (privately taken) action of an agent. Moreover, the principal has a deadline by which they need to complete all tasks. Using contracts to incentivize timely execution of tasks is prevalent in many time-sensitive tasks such as food delivery and contractors outsourcing tasks.
In Appendix B, we formally present the Knapsack Contracts model. We show that the Knapsack Contracts problem is reducible to the multi-choice stochastic knapsack with costs problem. Thus, providing approximation guarantees to the latter automatically translates to approximation guarantees for the former.
Proposition 1.1.
[Reduction to Multi-choice Stochastic Knapsack with Costs] Given an efficient algorithm that finds a policy that $c$-approximates the optimal policy for the MSKC problem, one can efficiently find a policy that $c$-approximates the optimal policy for the Knapsack Contracts problem.
1.1 Our Contributions and Techniques
In Section 2.1, we formally define the problem through which we explore the simplicity-optimality landscape when introducing costs, the Multi-choice Stochastic Knapsack with Costs problem (or MSKC). In this problem, each job is associated with multiple levels of execution, where each level has a different execution cost and induces a different size distribution. The decision-maker computes a policy to choose the (job, level) execution order, obtaining a utility of the value of jobs completed by the deadline minus the cost of all executed jobs. In Section 2.2 we formally define the different policies in our adaptiveness hierarchy.
As mentioned above, hard instances are ones where for some jobs, their cost almost “covers” their value. Thus, when executing such jobs, the policy is left with a small return on investment (ROI). In order to parameterize this, we inspect the Inverse of the ROI (or IOR) of a job (as the ROI diminishes, the IOR grows). We introduce an economically-meaningful parameter $\alpha$, which represents the worst IOR among all jobs—see Definition 2.1. We show this parameter is closely tied to the performance of non-adaptive policies in our problem. An overview of our main findings is given in Figure 1.
[…]
Figure 1: Hierarchy of policy adaptivity levels showing the approximation gaps between different classes.
1.1.1 $\Theta(\alpha)$ Adaptivity Gap.
In Section 3, we study the power of non-adaptive policies. Our results are the following.
Upper Bound.
We give an efficient algorithm for computing an ordered set of (job, execution level) pairs that give an $O(\alpha)$-approximation to the utility of the (intractable and highly adaptive) optimal policy (theorem 3.2).
@@ 1.1 Our Contributions and Techniques
Our algorithm formulates an LP for the MSKC problem where each pair $(i,j)$ has weight equal to the utility from executing job $i$ at execution level $j$, and size equal to the expected size of job $i$ at execution level $j$. In Dean et al. [10], a similar LP upper-bounds the utility of the optimal adaptive policy and can be approximated via the best non-adaptive policy. In our setting, the LP remains an upper bound on the optimal adaptive utility, but it may be a factor $\Omega(\alpha)$ away (see Section D.3). We show that some non-adaptive ordered set of job-level pairs achieves an $O(\alpha)$-approximation to the LP value.
We construct this set as follows. First, we efficiently obtain a well-structured LP solution in which every job is assigned integrally to at most one execution level, except possibly one job that may be split as a convex combination of two execution levels. We start from an arbitrary LP solution and shift weight along each job’s Pareto curve, iteratively making jobs integral while preserving the objective value. Given the resulting solution, we either apply the standard bang-per-buck ordering to the selected jobs, or take a single job if its expected utility exceeds a fixed threshold. The threshold is chosen to balance the potential loss when overflowing and the utility when selecting a high-utility job that is not part of the LP solution.
Lower Bound.
We show the existence of an instance such that any non-adaptive policy of the principal is bounded away by a factor of $\Omega(\alpha)$ in terms of utility from the optimal adaptive policy. This establishes a tight $\Theta(\alpha)$ adaptivity gap for the Knapsack Contracts problem (theorem 3.1). In fact, in the given instance, the gap exists between non-adaptive and stopping-adaptive policies.
1.1.2 A Constant Approximation with Limited Adaptivity
Motivated by the dependency of non-adaptive policies on ROI, in Section 4, we consider policies that go beyond non-adaptive policies, but still exhibit limited adaptivity. Our goal is to devise simple near-optimal policies, irrespective of the IOR parameter.
In Section 4.1 we first inspect whether policies that take a fixed, non-adaptive execution order, but can adaptively halt execution when future jobs are non-profitable, approximate the optimal utility. This is arguably the simplest class of policies that goes beyond non-adaptive policies, and can avoid incurring a high cost if the remaining time is insufficient. We show that this does not suffice; namely, we provide an instance where the ratio between the optimal utility and the utility of stopping-adaptive policies is $\Omega(\alpha^{1/3})$. In this instance, a good order of execution is obvious, but an optimal policy can skip over jobs that yield negative utility. This motivates us to study the next class of policies.
In Section 4.2 we give an efficient algorithm to compute a skipping-adaptive policy that gives a constant approximation to the utility of the optimal fully-adaptive policy. The policy is computed via a global time-indexed LP, an approach taken in [20, 24, 3], adapted to our setting.
1.1.3 Non-Stopping Policies
In Appendix F, we study an even more demanding setup, where the policy chooses an ordered set of items in advance, and incurs the cost of all of them, not only items that are inserted before the capacity is met. In the motivating example, this corresponds to a contractor who pre-commits to a fixed slate of subcontractors and must pay all agreed costs, even if earlier overruns make later jobs unprofitable or irrelevant. In the setting without costs (e.g., [10]), non-stopping and non-adaptive policies coincide. With costs, however, we show an $\Omega(\alpha)$ separation between them, and give an $O(\alpha)$-approximation non-stopping algorithm for the fully-adaptive optimum.
1.1.4 Information Gap
Our algorithm for computing a constant-approximation policy uses a substantial amount of distributional information, while our non-adaptive algorithm only uses the first moments of the size distributions and the overflow probability. In Appendix C, we establish an information gap: Consider any policy (adaptive or not) that has limited information about the jobs’ size distributions in the form of finitely many moments, as well as access to the probability of knapsack overflow for a single item. No such policy can guarantee an approximation factor better than $O(\alpha)$ with respect to the utility of the optimal policy with full access to the size distributions (theorem C.2). We note that the $O(\alpha)$ non-adaptive policy only uses the expectation (first moment information) and probability of overflow.
1.1.5 Relaxations.
In Appendix G, we study mild relaxations under which simpler policies achieve a constant-factor approximation. In Section G.2, we allow the IOR parameter to be arbitrarily large, but assume that for each job, the probability of consuming more than $1-\delta$ of the budget is small, for $\delta\in(0,1)$. We give an efficient algorithm that outputs a stopping-adaptive policy with an $O(1/\delta)$-approximation to the optimal fully-adaptive utility. We also show that non-adaptive policies fail to achieve an $\alpha$-independent approximation, even when job sizes are bounded by half the total time.
To prove this, we first study a variant where if a job completes after the deadline, the policy still receives its value and the process stops. For this variant, the algorithm from Section 3.2 gives a constant approximation to the LP value. We then run the same algorithm while treating the deadline as a $\delta$-fraction of the true deadline. In this view, bounded-size jobs correspond exactly to collecting the value of the first job that finishes after the (imagined) deadline, and shrinking the deadline costs only a factor $\delta$ in the objective.